# Phillip Long Computer Science & Engineering UC San Diego La Jolla, USA

### v1

@@ I Introduction
As a test bed, we use an unconditional multitrack music generation task to understand how metadata-based data filtering can affect downstream modeling, highlighting that the user rating data in PDMX can be used as a powerful measure for data quality and enable high quality subsets for training and fine-tuning generative models. Upon acceptance, PDMX will be released under a CC-BY license as a fully commercially viable dataset for symbolic music processing, as well as all related code for MusicRender and trained models.
@@ III Dataset
We created PDMX by scraping MuseScore [^1] https://musescore.com, a score-sharing online forum where community members (as well as formal music publishers) can upload their own sheet music arrangements and compositions as MusicXML files, including both licensed and public domain content. Besides the direct MusicXML files, we additionally scraped the metadata for each score in order to provide a cohesive understanding of the dataset, which includes tag-based descriptions of the score (i.e. composer, genre), community content in the form of ratings and comments, and most importantly the license for each score. We then selected for songs with specifically Public Domain Mark (i.e. already in the public domain prior to uploading) or CC-0 (i.e. explicitly released into the public domain) licenses. In total, our initial scrape yielded 254,077 MusicXML files, the largest dataset of MusicXML-like data to our knowledge, with a total 6,250.37 hours of music.

### v2

@@ I Introduction
As a test bed, we use an unconditional multitrack music generation task to understand how metadata-based data filtering can affect downstream modeling, highlighting that the user rating data in PDMX can be used as a powerful measure for data quality and enable high quality subsets for training and fine-tuning generative models. PDMX is released under a CC-BY license as a fully commercially viable dataset for symbolic music processing and available at our demo [^1] https://pnlong.github.io/PDMX.demo, as well as all related code for MusicRender and trained models.
@@ III Dataset
PDMX was created using data scraped from the online score-sharing forum MuseScore [^2] https://musescore.com [24], where community members (as well as formal music publishers) can upload their own sheet music arrangements and compositions as MusicXML files, including both licensed and public domain content. Besides the direct MusicXML files, we additionally scraped the metadata for each score in order to provide a cohesive understanding of the dataset, which includes tag-based descriptions of the score (i.e. composer, genre), community content in the form of ratings and comments, and most importantly the license for each score. We then selected for songs with specifically Public Domain Mark (i.e. already in the public domain prior to uploading) or CC-0 (i.e. explicitly released into the public domain) licenses. In total, our initial scrape yielded 254,077 MusicXML files, the largest dataset of MusicXML-like data to our knowledge, with a total 6,250.37 hours of music.
@@ V Conclusion
VI Acknowledgements
We thank Hao-Wen Dong for his efforts in scraping MuseScore and compiling the data used to create PDMX.